OpenStudy (anonymous):

The table below gives selected values for the function f(x). Use a trapezoidal estimation, with 6 trapezoids to approximate the value of the integral from 1 to 2 of f of x, dx. Give 3 decimal places for your answer.

OpenStudy (anonymous):

https://gyazo.com/0c664bbb486e2332ed456ba2ddaef650

OpenStudy (anonymous):

@terenzreignz Wanna see who can finish it first?

terenzreignz (terenzreignz):

I hate this kind of problem -_-

OpenStudy (anonymous):

Almost done anyways ;p

terenzreignz (terenzreignz):

OS was lagging on my end. But sure, you can have this one :)

OpenStudy (anonymous):

I was gonna win either way

terenzreignz (terenzreignz):

How typically like you to challenge me to unfair games :D

OpenStudy (anonymous):

I got 19.6

terenzreignz (terenzreignz):

Is it 7.4?

terenzreignz (terenzreignz):

Oh. My bad :D

OpenStudy (anonymous):

Did you just guess?

terenzreignz (terenzreignz):

No. I calculated.

OpenStudy (anonymous):

How'd you get 7.4?

terenzreignz (terenzreignz):

Trapezoidal rule. What's the correct answer?

OpenStudy (anonymous):

I don't know, I'm not given the answer.

OpenStudy (anonymous):

Leme cheat using my calculator skilz

OpenStudy (anonymous):

You were right

OpenStudy (anonymous):

God damn it

OpenStudy (anonymous):

roughly 7.36

OpenStudy (anonymous):

Either my method was wrong or I simply miscalculated.

terenzreignz (terenzreignz):

Maybe next time, you'll think twice before challenging me >:)

OpenStudy (anonymous):

I don't get how

terenzreignz (terenzreignz):

How did you know I was correct if you don't know how to do it?

OpenStudy (anonymous):

I did the same thing as always and still ending up with 19.6

terenzreignz (terenzreignz):

Why don't you show me what you did?

OpenStudy (anonymous):

My calculator

OpenStudy (anonymous):

Will a pic of my work do?

terenzreignz (terenzreignz):

sure.

OpenStudy (anonymous):

https://gyazo.com/b6b38482bd80bb2a7c46b00d994ba200

terenzreignz (terenzreignz):

Typos, galore. It seems this is the summation you did: \[\Large \sum_{n=1}^6 x_n\Delta f(x_n)\] Whereas this is what you should have done: \[\Large \sum_{n=1}^6 f(x_n)\Delta x_n\]

OpenStudy (anonymous):

Can you by any chance show me your work?

terenzreignz (terenzreignz):

No... it was done in my head :( But let me try to expound what you did was this \[\Large \left[\frac{\color{blue}{1+1.1}}{2}\right][\color{red}{3-1}]+\left[\frac{\color{blue}{1.1+1.3}}{2}\right][\color{red}{5-3}]...\] When what you should have done was this: \[\Large \left[\frac{\color{red}{3+1}}{2}\right][\color{blue}{1.1-1}]+\left[\frac{\color{red}{5+3}}{2}\right][\color{blue }{1.3-1.1}]...\]

OpenStudy (anonymous):

Ok, got it.